IPB

( Log In ) Log In is for TREC Teachers & Researchers only

 
Reply to this topicStart new topic
> July 28, 2006
Steve_Stevenoski
post Jul 30 2006, 02:27 PM
Post #1


Member
**

Group: TREC Team
Posts: 27
Joined: 19-July 06
Member No.: 39



July 28, 2006

We are still heading north and the ice looks like it is going to be more suitable for landing the helicopters. The ice is about 4 feet thick and startlingly blue against the white snow and the black of the deep ocean water. The ship has switched over to a 24-hour schedule as we prepare to do seismic work round the clock over the next four weeks.

Ice
IPB Image

Many little things had to be worked on today. Kevin and Matt were working to separate a bundle of cables and hoses that were used for single gun deployment. We needed the cables for the GI guns so all the tape holding the 120 plus feet of assorted hose and wire had to be cut off.

Hoses
IPB Image

By 9:30 Josh, a Coast Guard Welder was welding a shackle attachment point to the top of the cage for the crane cable. It is very windy and raw today. Icicles hang from the guns. After a half hour on the cage, Josh headed back to his bunk to put on a pair of long underwear. Kevin was on fire watch for about 2 hours as Josh welded. Someone has to watch constantly for fire caused by the sparks produced by the welder. After Josh was done welding for lunch break, I volunteered to remain on watch for a required 30 minutes after welding stops to make sure that there wasn't a stray ember that could start a fire left on the deck.

Welding the gun cage
IPB Image

After lunch, Matt went on fire watch as Josh finished up the welding.
Russell and I pulled the first of the GI gun cables. Dale worked on the splicing the remaining three GI gun cables. Larry discussed the time frame for completion of the cage for deployment with Dale. The plan for the evening was to do some coring beginning at about 10:00 or so, finishing about 6:00 am and transiting to the start of the first seismic area. The idea being that sometime tomorrow afternoon we would start firing the guns and collecting seismic data.

The weather has not cooperated with the helicopter operations. Every time we plan a flight, the weather turns bad, and we have to postpone. By 8:00 we had the second GI gun cable pulled and started to put tools away and clean up work areas for Saturday inspection. Dale and I cleaned toilets and swabbed the deck in the public toilets in the main lab, and Russell and I swept out the aft deck workroom.

The corers need to work on the fantail area right at the stern to deploy the corer. We had to move the protective tube and all the cable and hoses out of the way to open a direct path for the workroom on starboard to the stern. It was a four-person job. Marcy, Chase and Russell tugged and drag the mess to its new resting place. Captain Russell and Commander Jackson were making their rounds of the ship after we had moved the tube. We chatted about the ice and the future plans for the cruise and the weather.

Kelly was excited to get started with coring. She is a grad student at UAF, and has been looking for some science she can sink her hands into. If we get some good sediments she will be elbow deep in mud. She joined the coring team at about 10:00 to prepare all the coring gear for the first core of the cruise. They would start with a small core called a Benthos core to begin with.

Harm thought that we would begin helicopter operations tomorrow at 6:00. They were planning on making two or three flights to deploy anywhere from 3 to seven seismic instruments. If all continues to go well we should have the guns in the water by tomorrow evening Dale had just one last cable to splice.

11:50 - Still preparing the stern for taking the first core.

Crew
IPB Image
User is offlineProfile CardPM
Go to the top of the page
+Quote Post

Reply to this topicStart new topic
1 User(s) are reading this topic (1 Guests and 0 Anonymous Users)
0 Members:

 



- NSF Acknowledgment & Disclaimer Time is now: 23rd July 2016 - 01:13 PM
IPS Driver Error  

There appears to be an error with the database.
You can try to refresh the page by clicking here.

Error Returned

We apologise for any inconvenience